# EC-DiT: Scaling Diffusion Transformers with Adaptive Expert-Choice Routing

### v2

@@ 1 Introduction
The sparse Mixture-of-Experts (MoE) technique (Zoph et al., 2022; Lepikhin et al., 2020a; Du et al., 2022; Zhou et al., 2022) has proven to be an efficient method for scaling diffusion models like DiT. Sparse MoEs effectively scale up model capacity while maintaining relatively low computational overhead. In MoE architectures, the router selects a subset of parameters (experts) to process each group of token embeddings and then combines the final outputs from each expert. Current works (Yatharth Gupta, 2024; Fei et al., 2024; Xue et al., 2023) on sparse DiT often adopt a token-choice routing strategy (Zoph et al., 2022; Lepikhin et al., 2020a; Fedus et al., 2021), which routes each text token to the top-ranked experts. However, this routing strategy is suboptimal with diffusion models, which requires a more adaptive routing approach. Specifically, unlike autoregressive generation, DiT processes the entire image sequence at a time. This suggests that the global information of generated samples is accessible throughout the denoising procedure. Additionally, different image areas often contain varying levels of detail. Thus, an optimal router should adaptively allocate heterogeneous computation for different patterns.
Driven by these motivations, we propose scaling DiT with adaptive expert-choice routing (EC-DiT), which is of independent interest, and tailor it for text-to-image generation. This approach leverages global information from each image to optimize computational resource allocation that is aligned with varying image complexities, thus naturally aligning with the diffusion procedure. We scale EC-DiT up to 97 billion parameters with 64 experts. Figure 1 showcases the promising scalability of EC-DiT. Despite a less than 30% increase in computational overhead, our approach significantly improves training convergence, text-to-image alignment, and overall generation quality compared to dense and sparse variants with conventional token-choice MoEs (Lepikhin et al., 2020b; Du et al., 2022; Zoph et al., 2022). Notably, in text-to-image alignment evaluation, EC-DiT achieves a new GenEval score of 71.68% while maintaining competitive inference speed compared to other strong baselines. Our main contribution is summarized as follows:
@@ 2.1 Preliminaries
DiT processes images as sequences of patches in latent space. An input image is first equally divided into a grid of patches, then flattened and projected into a sequence of embeddings. We denote $\mathbf{x}\in\mathbb{R}^{S\times d_{x}}$ as the image sequence of length $S$ in latent space, with $d_{x}$ representing the model’s hidden dimension. Diffusion models progressively perform time-dependent denoising procedures on the image sequence $\mathbf{x}$. Given two observations $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$ and $\epsilon\sim p_{1}=\mathcal{N}\left(0,1\right)$, rectified flow (Liu et al., 2022; Lipman et al., 2022; Esser et al., 2024) defines a probability flow ordinary differential equation (ODE) to transfer between $p_{0}$ and $p_{1}$ with a straight-line path: $d\mathbf{x}_{t}=v_{\theta}(\mathbf{x}_{t},t)dt$. We parametrize the velocity (i.e., the path direction) $v_{\theta}(\mathbf{x}_{t},t)$ with DiT and optimize the following L2 objective (Liu et al., 2022; Esser et al., 2024):
$\mathcal{L}\left(\theta\right)=\mathbb{E}_{t\sim\pi\left(t\right),\epsilon\sim\mathcal{N}\left(0,1\right)}\left[\|(\mathbf{x}_{1}-\mathbf{x}_{0})-v_{\theta}(\mathbf{x}_{t},t)\|^{2}\right],\quad\mathbf{x}_{t}=t\mathbf{x}+(1-t)\epsilon,$ (1)
@@ 2.2 Text conditioning with cross-attention modules
In the text-to-image synthesis, the model generates image samples $\mathbf{x}$ conditioned on the text prompt $\mathbf{y}$. We utilize a text encoder $\text{Enc}\left(\mathbf{y}\right)\in\mathbb{R}^{L\times d_{y}}$ to extract the text embeddings of $\mathbf{y}$ where $L$ is the number of text tokens, and $d_{y}$ is the text’s hidden dimension. Following the approach in Chen et al. (2023), we incorporate a multi-head cross-attention layer (Vaswani et al., 2023) between the self-attention and FFN layers in each DiT block. The cross-attention $\mathbf{C}_{h}$ for head $h$ is written as:
@@ 2.3 Scaling DiT with expert-choice routing
Because the majority of computation takes place in the dense FeedForward (FFN) layers (Peebles & Xie, 2023) of DiT, replacing the FFN layer with a MoE layer is one efficient method of scaling. The introducing of the MoE layer decouples the effective computation from the model capacity. That is, we can effectively scale up the model capacity without significantly sacrificing the inference speed since only a selected subset of experts will be activated conditioned on the input representation. While this approach seems plausible and effective, it can be further improved for the following reasons:
@@
  • Load balancing. To improve the load balance of each expert, some sorts of auxiliary load balance loss (Shazeer et al., 2017; Lepikhin et al., 2020a; Du et al., 2022; Zoph et al., 2022) are often required for token-choice routing to work properly.
Therefore, we propose to scale up the diffusion models by better exploiting the structure of the diffusion process. Specifically, following Peebles & Xie (2023); Chen et al. (2023), we embed the timestep using a two-layer MLP with dimensionality equal to the hidden dimension $d_{x}$, and feed this embedding into each AdaLN layer. Additionally, as shown in Figure 2, let $\mathbf{x}_{s}$ be the output of the self-attention module, and $\mathbf{x}^{\prime}$ be the input to the router, i.e.,
@@
This affinity score tensor assesses the relevance between each pair of expert and input token.
Unlike the token-choice routing where each token $\mathbf{x}^{\prime}$ selects the top-$k$ experts from $\mathbf{A}_{s,i}$, EC-DiT works from the expert-choice (Zhou et al., 2022) view where each expert independently selects the top-$C$ tokens in the descending order from $\mathbf{A}_{s,i}$, and $C=B\times S\times f_{c}/E$ represents the average capacity of each expert, where $B$ is the batch size of image patches, $f_{c}$ denotes the capacity factor and reflects the average number of experts assigned to process each token.
@@
  • Adaptive computation. EC-DiT is implicitly aware of the multimodal context and denoising stages. As we will demonstrate in Section 3, EC-DiT effectively enables context-aware routing and dynamically adjusts its computation across both model depth and denoising timesteps.
  • Global context information. Diffusion models, unlike autoregressive models, process the entire sequence at each denoising step. As described in equation 6, EC-DiT leverages this global context, combining multi-modal information to optimize token-to-expert allocation.
@@ 3.1 Experiment Setup
Table 1: Model size and specifications across four backbone configurations. For Total Params., nE denotes EC-DiT with n experts in a single MoE layer. Activated Params. refer to the number of activated parameters in the forward pass. In Model Arch., #Head is the number of query heads, and #KV represents the number of key/value heads for grouped-query attention (GQA). *Note that a trade-off was made in the model depth for the M configuration, where Dense-M has 46 layers, while EC-DiT-M has 38 layers in order to be able to fit into the HBM of training accelerators.
@@
Following Chen et al. (2023), we adopt a modified DiT architecture with additional cross-attention modules for text-to-image generation. For all models evaluated, we use a 670M T5-based (Raffel et al., 2023) encoder for the text input and a 34M variational autoencoder (VAE) with 8 channels (CLIP-ViT-bigG) for the image input (Hessel et al., 2022). Both encoders remain frozen during DiT training. The transformer component is configured with four model sizes: XL, XXL, 3XL, and M, as detailed in Table 1. Starting from the second layer, we scale each dense model by interleavingly replacing all even-numbered dense layers with sparse MoE layers, where each expert is an FFN with dimensions identical to the dense counterpart. We set the capacity factor $f_{c}=2.0$ throughout the training and inference stages for all sparse models. We scale the XL-3XL dense models with 8, 16, and 32 experts. To further test the model’s scalability, we scale the M configuration with 64 experts. To explore the trade-off between model depth and width during scaling, we reduce the number of layers in EC-DiT-M-64E to 38, while maintaining 46 layers in the Dense. With these settings, our EC-DiT-3XL-32E reaches 32.15 billion parameters, and the largest EC-DiT-M-64E attains 97.21 billion parameters beyond all the current sparse diffusion models (Fei et al., 2024; Lin et al., 2024; Balaji et al., 2022).
@@
We collect and utilize approximately 1.2 billion text-image pairs from the Internet (McKinzie et al., 2024; Lai et al., 2023). The model resolution is set to $256\times 256$, with a patch size of 2. To accelerate training, we employ the masking technique proposed in (Zheng et al., 2023) with a masking ratio of 0.5. This results in the input sequence length of 128 per image. Model training is conducted on v4 and v5p TPUs with a batch size 4096. We use the RMSProp with momentum optimizer (Hinton, 2012) with a learning rate of 1e-4 and 20K warmup steps. All models are trained with Distributed Data Parallelism (DDP) or Fully Sharded Data Parallel (FSDP) for 800K steps. The expert dimension of EC-DiT is fully shared across the TPU mesh. To enhance training efficiency, we regroup the first two dimensions of the input sequence (batch_size, sequence_length) into (outer_batch, num_group, group_size). We maintain a group_size of 1024 for training and 512 for inference, which allows EC-DiT to process two images in a sequence at inference time.
@@
To assess the model’s text-to-image capabilities, we primarily use GenEval (Ghosh et al., 2023) and Davidsonian Scene Graph (DSG) scores (Cho et al., 2024). GenEval segments and detects objects in the generated samples to evaluate their alignment with the conditioned texts. For DSG, we use an internal version of Gemini to perform Visual Question Answering (VQA) to assess the text-image alignment. To evaluate the image quality of the generated images, we measure zero-shot Fréchet Inception Distance (FID) (Heusel et al., 2017) along with CLIP Score (Hessel et al., 2022) on the MS-COCO $256\times 256$ dataset using 30K samples (Lin et al., 2015). We also provide generated samples from a subset of Partiprompts (Yu et al., 2022) in Appendix C.
@@ 3.2 Text-to-image Alignment
Table 2: GenEval comparison. Our largest scaled model (EC-DiT-M with 64 experts) surpasses all existing models in overall performance and most sub-tasks of the GenEval Metrics (Ghosh et al., 2023), including the DPO-finetuned version of the current state-of-the-art, SD3 (Esser et al., 2024) (depth=38). The scores for models preceding ours are sourced from Esser et al. (2024). The best and second-best entries are highlighted.
@@
PixArt-$\alpha$ (Chen et al., 2023) 48.00 98.00 50.00 44.00 80.00 8.00 7.00
@@
SD3 (Esser et al., 2024) 68.00 98.00 84.00 66.00 74.00 40.00 43.00
SD3 (Esser et al., 2024) w/ DPO 71.00 98.00 89.00 73.00 83.00 34.00 47.00
@@
We present the comparative evaluation results on GenEval and DSG, both of which are commonly used to measure text-to-image alignment. Table 2 shows the comprehensive GenEval results of EC-DiT compared to current state-of-the-art models. Our largest sparse model, EC-DiT-M-64E, surpasses SD3 in the overall score and most sub-tasks. Notably, the solely pretrained EC-DiT-M-64E even outperforms the SD3 model fine-tuned with DPO. Moreover, our second-largest model, EC-DiT-3XL-32E, achieves comparable alignment capability to SD3 with DPO. It is worth noting that EC-DiT-M-32E is scaled from a 6B dense model, while the SD3 model listed in Table 2 is an 8B dense model with $512\times 512$ resolution. The difference in resolution may narrow the gap between our models and SD3, as SD3 generally yields higher GenEval scores at higher resolutions (Esser et al., 2024). This further underscores the potential benefits of scaling dense model with EC-DiT.
@@
We further validate the effectiveness of EC-DiT through DSG scores (Cho et al., 2024), as depicted in Figure 3. EC-DiT consistently improves DSG performance over the dense variants across different model configurations. Additionally, for each model size, increasing the number of experts further enhances text-image alignment. In both GenEval and DSG evaluations, increasing the model’s capacity with the EC-DiT directly enhances its text-following ability and leads to more accurately aligned image generation. A detailed DSG score breakdown is provided in Appendix A.

### v5

@@ 1 Introduction
The sparse Mixture-of-Experts (MoE) technique (Zoph et al., 2022; Lepikhin et al., 2020; Du et al., 2022; Zhou et al., 2022) has proven to be an efficient method for scaling diffusion models like DiT. Sparse MoEs effectively scale up model capacity while maintaining relatively low computational overhead. In MoE architectures, the router selects a subset of parameters (experts) to process each group of token embeddings and then combines the final outputs from each expert. Current works (Yatharth Gupta, 2024; Fei et al., 2024; Xue et al., 2023) on sparse DiT often adopt a token-choice routing strategy, which is initially proposed for scaling language models (Zoph et al., 2022; Lepikhin et al., 2020; Fedus et al., 2021). Despite technical differences, these methods, in principle, rely on routing each image token to the top-ranked experts. However, this routing strategy is suboptimal with diffusion models, which requires a more adaptive routing approach. Specifically, unlike autoregressive generation, DiT processes the entire image sequence at a time. This suggests that the global information of generated samples is accessible throughout the denoising procedure. Additionally, different image areas often contain varying levels of detail. Thus, an optimal router should adaptively allocate heterogeneous computation for different patterns.
Driven by these motivations, we propose scaling DiT with adaptive expert-choice routing (EC-DiT), which is of independent interest, and tailor it for text-to-image generation. This approach leverages global information from each image to optimize computational resource allocation that is aligned with varying image complexities, thus naturally aligning with the diffusion procedure. We scale EC-DiT up to 97 billion parameters with 64 experts. Figure 1 showcases the promising scalability of EC-DiT. Despite a less than 30% increase in computational overhead, our approach significantly improves training convergence, text-to-image alignment, and overall generation quality compared to dense and sparse variants with conventional token-choice MoEs (Yatharth Gupta, 2024; Fei et al., 2024). Notably, in text-to-image alignment evaluation, EC-DiT achieves a new GenEval score of 71.68% while maintaining competitive inference speed compared to other strong baselines. Our main contribution is summarized as follows:
@@ 2.1 Preliminaries
DiT processes images as sequences of patches in latent space. An input image is first equally divided into a grid of patches, then flattened and projected into a sequence of embeddings. We denote $\mathbf{x}\in\mathbb{R}^{S\times d_{x}}$ as the image sequence of length $S$ in latent space, with $d_{x}$ representing the model’s hidden dimension. Diffusion models progressively perform time-dependent denoising procedures on the image sequence $\mathbf{x}$. Given the observation $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$ from data distribution and noise $\mathbf{x}_{1}\sim p_{1}=\mathcal{N}\left(0,I\right)$, rectified flow (Liu et al., 2022; Lipman et al., 2022; Esser et al., 2024) transfers between two distributions $p_{0}$ and $p_{1}$ with a straight-line path $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$. This construction induces a probability flow ordinary differential equation (ODE) $d\mathbf{x}_{t}=v_{\theta}(\mathbf{x}_{t},t)dt$, where $v_{\theta}(\mathbf{x}_{t},t)$ represents the velocity direction of the path, parameterized by the DiT backbone. Since $v_{\theta}(\mathbf{x}_{t},t)=\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-\mathbf{x}_{0}$, we optimize the following L2 objective (Liu et al., 2022; Esser et al., 2024):
$\mathcal{L}\left(\theta\right)=\mathbb{E}_{t\sim\pi\left(t\right),\mathbf{x}_{1}\sim\mathcal{N}\left(0,I\right)}\left[\|(\mathbf{x}_{1}-\mathbf{x}_{0})-v_{\theta}(t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},t)\|^{2}\right],$ (1)
@@ 2.2 Text conditioning with cross-attention modules
In the text-to-image synthesis, the model generates image samples $\mathbf{x}$ conditioned on the text prompt $\mathbf{y}$. We utilize a text encoder $\text{Enc}\left(\mathbf{y}\right)\in\mathbb{R}^{L\times d_{y}}$ to extract the text embeddings of $\mathbf{y}$ where $L$ is the number of text tokens, and $d_{y}$ is the text’s hidden dimension. Following the approach in Chen et al. (2023a), we incorporate a multi-head cross-attention layer (Vaswani et al., 2023) between the self-attention and FFN layers in each DiT block. The cross-attention $\mathbf{C}_{h}$ for head $h$ is written as:
@@ 2.3 Scaling DiT with expert-choice routing
Because the majority of computation takes place in the dense FeedForward (FFN) layers (Peebles & Xie, 2023) of DiT, replacing the FFN layer with an MoE layer is one efficient method of scaling. Introducing the MoE layer decouples the effective computation from the model capacity. That is, we can effectively scale up the model capacity without significantly sacrificing the inference speed since only a selected subset of experts will be activated conditioned on the input representation. While this approach seems plausible and effective, it can be further improved for the following reasons:
@@
  • Load balancing. To improve the load balance of each expert, some sorts of auxiliary load balance loss (Shazeer et al., 2017; Lepikhin et al., 2020; Du et al., 2022; Zoph et al., 2022) are often required for token-choice routing to work properly.
Therefore, we propose to scale up the diffusion models by better exploiting the structure of the diffusion process. Specifically, following Peebles & Xie (2023); Chen et al. (2023a), we embed the timestep using a two-layer MLP with dimensionality equal to the hidden dimension $d_{x}$, and feed this embedding into each AdaLN layer. Additionally, as shown in Figure 2, let $\mathbf{x}_{s}$ be the output of the self-attention module, and $\mathbf{x}^{\prime}$ be the input to the router, i.e.,
@@
This affinity score tensor assesses the relevance between each pair of expert and input token. Unlike the token-choice routing where each token $\mathbf{x}^{\prime}$ selects the top-$k$ experts from $\mathbf{A}_{s,i}$, EC-DiT works from the expert-choice (Zhou et al., 2022) view where each expert independently selects the top-$C$ tokens in the descending order from $\mathbf{A}_{s,i}$, and $C=S\times f_{c}/E$ represents the average capacity of each expert, where $f_{c}$ denotes the capacity factor and reflects the average number of experts assigned to process each token.
@@
  • Adaptive computation. EC-DiT is implicitly aware of the textual context and denoising stages. As we will demonstrate in Section 3, EC-DiT effectively enables context-aware routing and dynamically adjusts its computation across both model depth and denoising timesteps.
  • Global context information. Diffusion models, unlike autoregressive models, process the entire sequence at each denoising step. As described in equation 6, EC-DiT leverages this global context, combining textual and image information to optimize token-to-expert allocation.
@@ 3.1 Experiment Setup
Table 1: Model size and specifications across four backbone configurations. For Total Params., nE denotes EC-DiT with n experts in a single MoE layer. Activated Params. refer to the average number of activated parameters per token in the forward pass. The detailed calculation is depicted in Appendix D. In Model Arch., #Head is the number of query heads, and #KV represents the number of key/value heads for grouped-query attention (GQA). *Note that a trade-off was made in the model depth for the M configuration, where Dense-M has 46 layers, while EC-DiT-M has 38 layers in order to be able to fit into the HBM of training accelerators.
@@
Following Chen et al. (2023a), we adopt a modified DiT architecture with additional cross-attention modules for text-to-image generation. For all models evaluated, we use a 670M clip-based text encoder with the T5 tokenizer (Raffel et al., 2023) for the text input and a 34M variational autoencoder (VAE) with 8 channels (CLIP-ViT-bigG) for the image input (Hessel et al., 2022). Both encoders remain frozen during DiT training. The transformer component is configured with four model sizes: XL, XXL, 3XL, and M, as detailed in Table 1. Starting from the second layer, we scale each dense model by interleavingly replacing all even-numbered dense layers with sparse MoE layers, where each expert is an FFN with dimensions identical to the dense counterpart. We set the capacity factor $f_{c}=2.0$ throughout the training and inference stages for all sparse models. We scale the XL-3XL dense models with 8, 16, and 32 experts. To further test the model’s scalability, we scale the M configuration with 64 experts. To explore the trade-off between model depth and width during scaling, we reduce the number of layers in EC-DiT-M-64E to 38, while maintaining 46 layers in the Dense. With these settings, our EC-DiT-3XL-32E reaches 32.15 billion parameters, and the largest EC-DiT-M-64E attains 97.21 billion parameters beyond all the current sparse diffusion models (Fei et al., 2024; Lin et al., 2024; Balaji et al., 2022).
@@
We collect and utilize approximately 1.2 billion text-image pairs from the Internet (McKinzie et al., 2024; Lai et al., 2023). The model resolution is set to $256\times 256$, with a patch size of $2\times 2$. To accelerate training, we employ the masking technique proposed in (Zheng et al., 2023) with a masking ratio of 0.5. This results in the input sequence length of 128 per image. Model training is conducted on v4 and v5p TPUs with a batch size 4096. We use the RMSProp with momentum optimizer (Hinton, 2012) with a learning rate of 1e-4 and 20K warmup steps. All models are trained with Distributed Data Parallelism (DDP) or Fully Sharded Data Parallel (FSDP) for 800K steps. The expert dimension of EC-DiT is fully shared across the TPU mesh. To enhance training efficiency, we regroup the first two dimensions of the input sequence (batch_size, sequence_length) into (outer_batch, num_group, group_size). We maintain a group_size of 1024 for training and 512 for inference, which allows EC-DiT to process two images in a sequence at inference time.
@@
To assess the model’s text-to-image capabilities, we primarily use GenEval (Ghosh et al., 2023) and Davidsonian Scene Graph (DSG) scores (Cho et al., 2024). GenEval segments and detects objects in the generated samples to evaluate their alignment with the conditioned texts. For DSG, we use an internal version of Gemini to perform Visual Question Answering (VQA) to assess the text-image alignment. To evaluate the image quality of the generated images, we measure zero-shot Fréchet Inception Distance (FID) (Heusel et al., 2017) along with CLIP Score (Hessel et al., 2022) on the MS-COCO $256\times 256$ dataset using 30K samples (Lin et al., 2015). We also provide generated samples from a subset of Partiprompts (Yu et al., 2022) in Appendix E.
@@ 3.2 Text-to-image Alignment
Table 2: GenEval comparison. Our largest scaled model (EC-DiT-M with 64 experts) surpasses all existing models in overall performance and most sub-tasks of the GenEval Metrics (Ghosh et al., 2023), including the DPO-finetuned version of the current state-of-the-art, SD3-Large (Esser et al., 2024) (depth=38, resolution $512\times 512$). The scores for models preceding ours are sourced from Esser et al. (2024). The best and second-best entries are highlighted.
@@
PixArt-$\alpha$ (Chen et al., 2023a) 48.00 98.00 50.00 44.00 80.00 8.00 7.00
@@
SD3-Large (Esser et al., 2024) 68.00 98.00 84.00 66.00 74.00 40.00 43.00
SD3-Large (Esser et al., 2024) w/ DPO 71.00 98.00 89.00 73.00 83.00 34.00 47.00
@@
We present the comparative evaluation results on GenEval and DSG, both of which are commonly used to measure text-to-image alignment. Table 2 shows the comprehensive GenEval results of EC-DiT compared to current state-of-the-art models. Specifically, our largest sparse model, EC-DiT-M-64E (resolution $256\times 256$), reaches GenEval of 71.68% and surpasses SD3 (resolution $512\times 512$) in the overall score (68.00%) and most sub-tasks. Notably, the solely pretrained EC-DiT-M-64E without fine-tuning stage even outperforms the SD3 model fine-tuned with DPO. Notably, both DPO fine-tuning and training with higher resolution positively impact the GenEval score, as SD3 generally yields higher GenEval scores equipped with these techniques (Esser et al., 2024). This further underscores the potential benefits of scaling dense model with EC-DiT. Furthermore, EC-DiT-3XL-32E, with 5.18B activated parameters, achieves a nearly equivalent GenEval score of 70.91% compared to the 71.00% score from SD3-Large with DPO, with only 64% activation size of the latter. The EC-DiT-3XL-32E’s inference time is only 33.8% of the 8B Dense-M. These results demonstrate that our scaling approach can achieve performance comparable to significantly larger dense models, even without additional fine-tuning.
@@
We further validate the effectiveness of EC-DiT through DSG scores (Cho et al., 2024), as depicted in Figure 3. EC-DiT consistently improves DSG performance over the dense variants across different model configurations. Additionally, for each model size, increasing the number of experts further enhances text-image alignment. In both GenEval and DSG evaluations, increasing the model’s capacity with the EC-DiT directly enhances its text-following ability and leads to more accurately aligned image generation. A detailed DSG score breakdown is provided in Appendix B.